Gucchi:

The length of the hypotenuse of a right triangle is 157 units. The length of one leg of the triangle is 132 units. Lara wrote the following step to find the length of the unknown leg: Length of the unknown leg = 157^2 − 132^2 = 24,649 − 17,424 = 7,225 units Which statement best explains whether Lara's step is correct or incorrect? It is correct because the length of the unknown side is the difference of the lengths of the sides. It is correct because the length of the unknown side is the difference of the squares of the sides. It is incorrect because the length of the unknown side is the square root of 42,073. It is incorrect because the length of the unknown side is the square root of 7,225.

Gucchi:

Am I right with the last one?

Livxxx:

Yes you are correct

Gucchi:

explanation?

Gucchi:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @yourgrandpa a^2 + b^2 = c^2 132^2 + b^2 = 157^2 17424 + b^2 = 24,649 b^2 = 24,649 - 17424 b^2 = 7225 b = square root 7225 b = 85 <== the missing leg is 85 units Lauras step is correct...she rearranged the formula making it : b^2 = c^2 - a^2.....and subbed in numbers correctly. She just didn't finish \(\color{#0cbb34}{\text{End of Quote}}\) thats copy and paste from brainly

pooja195:

a^2+b^2=c^2 We know: 132^2+b^2=157^2 We we are missing b^2 as is To solve this we need to rearrange the formula

Gucchi:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @pooja195 a^2+b^2=c^2 We know: 132^2+b^2=157^2 We we are missing b^2 as is To solve this we need to rearrange the formula \(\color{#0cbb34}{\text{End of Quote}}\) I see.

Gucchi:

?

pooja195:

b=sqrt c^2-a^2

pooja195:

So you will do: 157^2-132^2=?

Gucchi:

24,649 - 17, 424?

pooja195:

Yes

pooja195:

What do you get?

Gucchi:

Answer: 7,225

Gucchi:

So It is D?

Gucchi:

I get it now.

pooja195:

Yes, because she doesn't mention a squareroot she only mentions the number which is why it is wrong so D is right It is incorrect because the length of the unknown side is the square root of 7,225.

Gucchi:

Alright, makes sense now. Thanks

pooja195:

You're welcome :-)